OpenStudy (shelby1290):

The number of bees in a hive is 1000 on June 1 and doubles every month. This can be expressed as N=1000 X 2^t, where N represents the number of bees and t represents time, in months. a) Find the number of bees after 2,3,4, and 5 months b) What does t=0 represent in this situation? c)Is it possible for t to be -1? What does this mean? d) When were there 125 bees? Explain.

OpenStudy (zale101):

N=1000 x 2^t x is not another variable, it's multiplication?

OpenStudy (shelby1290):

yes it's multiplication

OpenStudy (zale101):

have you tried doing part a at least?

OpenStudy (shelby1290):

I just have to sub 2,3,4 and 5 for t right?

OpenStudy (zale101):

Yes

OpenStudy (zale101):

To get N which represents the number of bees

OpenStudy (shelby1290):

1000x2^2=4000 in 2months 1000x2^3=8000 in 3 months 1000x2^4=16,000 in 4 months 10000x2^5=32,000 in 5 months

OpenStudy (shelby1290):

I don't really understand part b

OpenStudy (shelby1290):

@Zale101 ^

OpenStudy (zale101):

b) What does t=0 represent in this situation? What happens to the function when t=0?

OpenStudy (shelby1290):

1000 x 2^0=1000 the number of bees stays the same when t is 0? idk

OpenStudy (zale101):

In your question, it says "The number of bees in a hive is 1000 on June 1 " So, when t=0, the number of bees is 1000. That is the same number of bees starting on June 1.

OpenStudy (shelby1290):

ohh i see

OpenStudy (shelby1290):

for part c... 1000 X 2^-1=500

OpenStudy (shelby1290):

Are those the number of bees before June 1st?

OpenStudy (zale101):

Yes.

OpenStudy (zale101):

For part D, all you have to do is sub 125 to n and solve for t using logs.

OpenStudy (shelby1290):

okay so 125=1000 x2^t but what are logs?

OpenStudy (zale101):

These are logs https://www.mathsisfun.com/algebra/logarithms.html

OpenStudy (zale101):

\( log_{10}(\frac{125}{1000})=log_{10} 2^t\)

OpenStudy (shelby1290):

I've never learned about logs before based on the example would be it be 5x5x5=125 i'm just guessing

OpenStudy (zale101):

125/1000 is reduced to 1/8 \(log_{10}(\frac{1}{8}=log_{10}2^t\)

OpenStudy (zale101):

well, try to do further study on logarithms because you might get tested on it.

OpenStudy (shelby1290):

@Zale101 so would my answer be log10(1/8=log10 2t

OpenStudy (shelby1290):

or..

OpenStudy (shelby1290):

something else

OpenStudy (zale101):

\((log_{10}(1/8))= (log_{10}(2^t))\) \({log_{10}\frac{1}{8}}=t~log_{10}2\) \(\LARGE \frac{(log_{10}(\frac{1}{8})}{(log_{10}(2))}=\large t\)

OpenStudy (zale101):

what would that equal to?

OpenStudy (shelby1290):

somehow i got 1/4

OpenStudy (zale101):

http://www.chilimath.com/algebra/advanced/log/images/466x428xrules,P20of,P20exponents.gif.pagespeed.ic.1z7-Ekl5L5.png Here are some log rules for you to know and learn of how i solved it.

OpenStudy (zale101):

The answer is not 1/4

OpenStudy (zale101):

http://www.wolframalpha.com/input/?i=%28log_10%281%2F8%29%29%2F%28log_10%282%29%29%3Dt You can just plug it to the calculator or wolfram.

OpenStudy (shelby1290):

@Zale101 alright so the final answer is -3?